OpenStudy (anonymous):

upper limit is 1 lower limit is 0 equation is x^2dx/x^3+4x^2+5x+2=??

OpenStudy (anonymous):

\[\left( \int\limits_{0}^{1} x ^{2}dx \right)/x^3+4x^2+5x+2= \] means nothing…

OpenStudy (anonymous):

just u sub the denominator and the x^2 will cancel or do partial fractions